OpenStudy (anonymous):

Find the product: (4x-5y)^2= (4x-5y)(4x-5y)= (4x)(4x)=16x^2 (4x)(-5y)=-20xy (-5y)(4x)=-20xy (-5y)(-5y)=25y^2 16x^2-20xy-20xy+25y^2= 16x^2+25y^2 Is this correct?

OpenStudy (phi):

This part is good 16x^2-20xy-20xy+25y^2 but how did you make the -20xy - 20xy go away! you can combine them

OpenStudy (phi):

add "like terms" like terms have the same variables (letters). Just add their coefficients -20xy - 20xy = -40xy

OpenStudy (anonymous):

ok, so I originally had 16x^2-40xy + 25y^2, doubted myself

OpenStudy (phi):

you check yourself by replacing x and y with numbers. for example, let x=1 and y=1 (makes it easy) (4x-5y)^2 becomes (4-5)^2= -1*-1= 1 now try 16x^2-40xy + 25y^2. It becomes 16*1*1-40*1*1+25*1*1 or 16-40+25= 1 on the other hand 16x^2+25y^2 is 16+25= 41 which is definitely not 1

OpenStudy (anonymous):

ok, thank you for explaining that. it definitely helps!!:)